Calculate the reflection coefficient R2 of the output coupler of L cm cavity tube of a laser of wavelength is 632.8nm which has the threshold gain factor and loss factor of 0.15 m-1 and 0.05m-1 respectively with one mirror of R1% reflecting.
L = 0.91 R1% = 98.62
1
Expert's answer
2020-05-21T13:09:25-0400

Assume that laser operates in threshold condition "G_L=(G_L)_{th}=1" , where "G_L" is the Loop Gain.


"(G_L)_{th}=1=R_1\cdot R_2\cdot exp(2(\beta-\alpha)L)".


In our case "R_1=0.9862" , "\alpha=0.05m^{-1}" , "\beta=0.15m^{-1}" and "L=0.91m"


So,


"R_2=\frac{1}{R_1\cdot exp(2(\beta-\alpha)L)}=\frac{1}{0.9862\cdot exp(2\cdot(0.15-0.05)\cdot0.91)}=0.8453" Answer